# The Simons Observatory: Science Goals and Forecasts for the Enhanced Large Aperture Telescope

### v2

@@ 3.6 Insights into the polarized Galactic interstellar medium
Arcminute-resolution millimeter-wave maps of the SO footprint, which includes much of the Galactic plane, will be valuable for answering many questions about the physical mechanisms responsible for polarized Galactic emission, via measurements of the dust and synchrotron spectral energy distributions, and detections or upper limits on polarized CO line emission and anomalous microwave emission (Hensley et al. 2022). Together with large-area starlight polarization surveys such as PASIPHAE (Tassis et al. 2018), these data will yield new insights into the physics of Galactic emission and the multi-scale structure of the magnetized interstellar medium (Hensley et al. 2019; Fissel et al. 2019).
@@ 3.10 The unexplored millimeter transient sky
Recent years have seen new opportunities for millimeter time-domain astronomy. One reason is that the discovery of young ( days-old) transients by optical surveys has become routine, enabling fast turnaround follow-up observations using sensitive millimeter interferometers (Ho et al. 2019; Maeda et al. 2021; Andreoni et al. 2022; Berger et al. 2023). In addition, transients have been discovered blindly at millimeter wavelengths for the first time, in CMB survey data (Whitehorn et al. 2016; Guns et al. 2021; Naess et al. 2021a; Li et al. 2023; Biermann et al. 2024), and these data have also been used to put millimeter-wavelength flux limits on known extragalactic transients (Hervías-Caimapo et al. 2024b).
@@
Projected detection rates are presented in Table 4. For fast events (durations between 1 and 10 days), which would include some LGRB reverse shocks and on-axis LGRB afterglows, we use the upper limit on the rate computed by SPT (Guns et al. in prep). For the on-axis LGRB rate we use Lien et al. (2014). For the LGRB and TDE luminosities at 145 GHz we use theoretical predictions (Metzger et al. 2015; Eftekhari et al. 2022) and caution that these are uncertain given the small number of observed events. We take the LLGRB rate from Soderberg et al. (2006), and for the luminosity we use the observed values of GRB 980425/SN 1998bw (Kulkarni et al. 1998) and GRB 171205A/SN 2017iuk (Perley et al. 2017). For AT2018cow-like FBOTs we use observed 100 GHz luminosities (Ho et al. 2022, 2023b) and the volumetric rate from Ho et al. (2023a). For ordinary SNe we use an observationally constrained rate (Li et al. 2011; Perley et al. 2020), and luminosities of the handful of events detected at millimeter wavelengths to date (Berger et al. 2023). We find that ordinary SNe would likely only be detected out to a few Mpc, and thus we do not list these in Table 4, as we expect the detection rate for extragalactic normal SNe to be low. For interacting SNe we use a volumetric rate from ZTF (Perley et al. 2020) and a range of luminosities from the modeling in Yurk et al. (2022).
@@ 4 Summary
The science goals and forecasts for the wide-area survey conducted with the fully-populated LATR are described in Sec. 3 and summarized in Table 2. These include improved constraints on the scale dependence and Gaussianity of the primordial perturbations (reaching ); improved constraints on new light relativistic species (reaching ); tight constraints on the sum of the neutrino masses ( eV, or 0.015 eV in combination with LiteBIRD); percent-level constraints on the amplitude of density fluctuations at redshifts via CMB lensing cross-correlations with LSS surveys; and a tSZ-selected galaxy cluster sample comprising 33,000 objects, including at . In general, the high-SNR component-separated blackbody temperature and Compton- maps, as well as the reconstructed gravitational lensing maps, from the complete SO LAT survey will enable a broad range of cosmological and astrophysical science, which we have only briefly covered in this paper. As evidenced by progress in cosmology in recent decades, a data set this rich will enable additional science that has yet to be envisioned or forecast here.
@@ Appendix A Sensitivity and forecasting assumptions
Table 1 provides map-domain sensitivities for the full-depth maps expected at the conclusion of the SO LAT survey in 2034. To obtain these numbers, we take a similar approach to that in Ade et al. (2019), using an atmospheric noise model informed by ACT observations combined with detector NET sensitivities estimated using bolocalc (Hill et al. 2018). For the nominal-survey SO forecasts described in Ade et al. (2019), it was assumed that 20% of the data collected over a 5-year duration would be usable for cosmological analysis (after accounting for uptime, CMB field availability, and data quality cuts, all of which were estimated based on historical data from ACT). For this work, we use the same detector NET sensitivities and 20% observation efficiency, but make the following changes. Rather than assuming a 5-year survey with the OT configuration from Ade et al. (2019) (comprising 1 LF, 4 MF, and 2 UHF OTs), we assume 3 years of nominal SO observations with that tube configuration, followed by 6 years of observations with the fully populated LATR (comprising 1 LF, 8 MF, and 4 UHF OTs, as described in Sec. 2). Ade et al. (2019) assumed that the sky fraction observed would comprise , with a post-masking footprint of available for cosmological analysis (avoiding regions of Galactic contamination and poor cross-linking near the footprint edges). In this work, following updates to the SO scan strategy that accommodate a wider range of science goals, we assume that throughout the 9-year survey duration the SO LAT will map a sky fraction . These assumptions lead to the full-depth map sensitivities given in Table 1. Further details for the cosmological science forecasts and the point-source and transient forecasts are given below.
@@
Transient and point source science: We use two types of sensitivities for our transient and point source forecasts. First, we estimate our sensitivities to point sources from a single observation, that is, for a point on the sky that drifts once through the entire focal plane, which is scanning in azimuth at a fixed elevation. We base these sensitivities on simulated maps made with our nominal observing strategy, in which the majority of observations occur at elevation (70.6% of observations), with a minority at (15.7%) and at (13.7%). We use a fixed set of detector NETs for these simulations, and we then rescale the maps to correspond to the NETs used for the full-depth cosmological forecasts listed in Table 1. We take the median map depth, in units of , for each of the observing elevations, and calculate their weighted mean based on the observing fraction at each elevation. Second, we estimate our sensitivities to point sources in the final, full-survey maps by spreading the array sensitivity accumulated over the survey duration over the survey area, using and an observing efficiency of , as described above. For both single-observation and full-depth maps, we obtain point source sensitivities by converting the map depth in to mJy using the beam FWHMs listed in Table 1 and applying an inverse-variance weighting with an atmospheric noise spectrum with a power-law index of and for GHz, respectively. Here, the values are estimated based on ACT data.131313Note that Ade et al. (2019) also made use of a set of values for the noise model used in forecasts (which is used in the cosmological forecasting in this paper as well), but there the term would have been better called a reference multipole, , rather than a knee multipole, which is usually defined as the multipole where the red noise caused by the atmosphere reaches the same level as the white noise; Ade et al. (2019) instead defined a “red noise” amplitude, , to scale the reference multipole. The relationship between the two is: , where is the power-law index of the red noise. For the point source forecasts in this paper, our choices of and are independent of the model used in Ade et al. (2019), but are still chosen to be consistent with the atmospheric noise measured by ACT. For the full-depth maps, we include additional “noise” contributions from the CMB and the CIB, since these terms dominate over the atmospheric noise in this case (other background sources like unresolved AGN are ignored). We ignore these terms for the single-observation maps since they are subdominant. The resulting sensitivities for both the single-observation and full-survey cases are given in Table 3.

### v3

@@ 3.6 Insights into the polarized Galactic interstellar medium
Arcminute-resolution millimeter-wave maps of the SO footprint, which includes much of the Galactic plane, will be valuable for answering many questions about the physical mechanisms responsible for polarized Galactic emission, via measurements of the dust and synchrotron spectral energy distributions across the sky, and detections or upper limits on polarized CO line emission and anomalous microwave emission (AME; Hensley et al. 2022). SO will detect AME polarization if the AME is polarized at the 0.1% level, which is sufficient to distinguish between competing theories of ultrasmall grain alignment (Hensley et al. 2022), such as resonance paramagnetic alignment (Lazarian & Draine 2000; Hoang et al. 2013), or suppression of alignment due to quantum mechanical effects (Draine & Hensley 2016). Additionally, SO will determine what fraction of the AME could be contributed by magnetic dipole emission from ferromagnetic grains, which are distinguished by their unique polarization signature (Draine & Hensley 2013). See Dickinson et al. (2018) for a recent review. Together with large-area starlight polarization surveys such as PASIPHAE (Tassis et al. 2018), these data will yield new insights into the physics of Galactic emission and the multi-scale structure of the magnetized interstellar medium (Hensley et al. 2019; Fissel et al. 2019).
@@ 3.10 The unexplored millimeter transient sky
Recent years have seen new opportunities for millimeter time-domain astronomy. One reason is that the discovery of young ( days-old) transients by optical surveys has become routine, enabling fast turnaround follow-up observations using sensitive millimeter interferometers (Ho et al. 2019; Maeda et al. 2021; Andreoni et al. 2022; Berger et al. 2023). In addition, transients have been discovered blindly at millimeter wavelengths for the first time, in CMB survey data (Whitehorn et al. 2016; Guns et al. 2021; Naess et al. 2021a; Li et al. 2023; Biermann et al. 2024), and these data have also been used to put millimeter-wavelength flux limits on known extragalactic transients (Herv´ıas-Caimapo et al. 2024).
@@
Projected detection rates are presented in Table 4. For fast events (durations between 1 and 10 days), which would include some LGRB reverse shocks and on-axis LGRB afterglows, we use the upper limit on the rate computed by SPT (Guns et al. in prep). For the on-axis LGRB rate we use Lien et al. (2014). For the LGRB and TDE luminosities at 145 GHz we use theoretical predictions (Metzger et al. 2015; Eftekhari et al. 2022) and caution that these are uncertain given the small number of observed events. We take the LLGRB rate from Soderberg et al. (2006), and for the luminosity we use the observed values of GRB 980425/SN 1998bw (Kulkarni et al. 1998) and GRB 171205A/SN 2017iuk (Perley et al. 2017). For AT2018cow-like FBOTs we use observed 100 GHz luminosities (Ho et al. 2022, 2023b) and the volumetric rate from Ho et al. (2023a). For ordinary SNe we use an observationally constrained rate (Li et al. 2011; Perley et al. 2020), and luminosities of the handful of events detected at millimeter wavelengths to date (Berger et al. 2023). We find that ordinary SNe would likely only be detected out to a few Mpc, and thus we do not list these in Table 4, as we expect the detection rate for extragalactic normal SNe to be low. For interacting SNe we use a volumetric rate from the Zwicky Transient Facility (Perley et al. 2020) and a range of luminosities from the modeling in Yurk et al. (2022).
@@ 4 Summary
The science goals and forecasts for the wide-area survey conducted with the fully-populated LATR are described in Sec. 3 and summarized in Table 2. The forecasts presented in this work include only statistical errors; our understanding of systematic errors will be refined in future work, particularly as the instrument is now taking data. These include improved constraints on the scale dependence and Gaussianity of the primordial perturbations (reaching ); improved constraints on new light relativistic species (reaching ); tight constraints on the sum of the neutrino masses ( eV, or 0.015 eV in combination with LiteBIRD); percent-level constraints on the amplitude of density fluctuations at redshifts via CMB lensing cross-correlations with LSS surveys; and a tSZ-selected galaxy cluster sample comprising 33,000 objects, including at . In general, the high-SNR component-separated blackbody temperature and Compton- maps, as well as the reconstructed gravitational lensing maps, from the complete SO LAT survey will enable a broad range of cosmological and astrophysical science, which we have only briefly covered in this paper. As evidenced by progress in cosmology in recent decades, a data set this rich will enable additional science that has yet to be envisioned or forecast here.
@@ Appendix A Sensitivity and forecasting assumptions
Table 1 provides map-domain sensitivities for the full-depth maps expected at the conclusion of the SO LAT survey in 2034. To obtain these numbers, we take a similar approach to that in Ade et al. (2019), using an atmospheric noise model informed by ACT observations combined with detector NET sensitivities estimated using bolocalc (Hill et al. 2018). For the nominal-survey SO forecasts described in Ade et al. (2019), it was assumed that 20% of the data collected over a 5-year duration would be usable for cosmological analysis (after accounting for uptime, CMB field availability, and data quality cuts, all of which were estimated based on historical data from ACT). For this work, we use the same detector NET sensitivities and 20% observation efficiency, but make the following changes. Rather than assuming a 5-year survey with the OT configuration from Ade et al. (2019) (comprising 1 LF, 4 MF, and 2 UHF OTs), we assume 3 years of nominal SO observations with that tube configuration, followed by 6 years of observations with the fully populated LATR (comprising 1 LF, 8 MF, and 4 UHF OTs, as described in Sec. 2). Ade et al. (2019) assumed that the sky fraction observed would comprise , with a post-masking footprint of available for cosmological analysis (avoiding regions of Galactic contamination and poor cross-linking near the footprint edges). In this work, following updates to the SO scan strategy that accommodate a wider range of science goals, we assume that throughout the 9-year survey duration the SO LAT will map a sky fraction . These assumptions lead to the full-depth map sensitivities given in Table 1. For completeness, we provide the “goal” and “baseline” per-OT NET sensitivities in Table 5. Further details for the cosmological science forecasts and the point-source and transient forecasts are given below.
| Band | LF1 | LF2 | MF1 | MF2 | UHF1 | UHF2 |
| Frequency [GHz] | 27 | 39 | 93 | 145 | 225 | 280 |
| Baseline NET [K] | 48 | 24 | 10.8 | 13.4 | 21.2 | 50.9 |
| Goal NET [K] | 35 | 18 | 7.8 | 8.4 | 14.1 | 35.4 |
Forecasts in this work and in Ade et al. (2019) are based on “baseline” and “goal” noise models. This table provides the associated NET per OT; note that each OT includes three 150-mm detector arrays. The baseline model represents the requirements for SO, while goal represents the expected performance. These models include parameters to encapsulate the key drivers for instrumental sensitivity, including: the emission from and opacity of the atmosphere; the properties of the telescope (emission, spill, etc.); properties of the receiver and cryogenic optics (emission, losses, and reflections from the windows, lenses, and cryogenic cold stop); properties of the filters that define the passbands; and detector properties (yield, noise, etc.). The sensitivities presented here were computed at the start of the design process, when many instrumental details were yet to be finalized. For this reason, the parameter choices are uninformative and are not reported here. Pre-deployment testing of the MF OTs indicates performance exceeding the baseline requirement and consistent with goal performance (Sierra et al. 2025).
@@
Transient and point source science: We use two types of sensitivities for our transient and point source forecasts. First, we estimate our sensitivities to point sources from a single observation, that is, for a point on the sky that drifts once through the entire focal plane, which is scanning in azimuth at a fixed elevation. We base these sensitivities on simulated maps made with our nominal observing strategy, in which the majority of observations occur at elevation (70.6% of observations), with a minority at (15.7%) and at (13.7%). We use a fixed set of detector NETs for these simulations, and we then rescale the maps to correspond to the NETs used for the full-depth cosmological forecasts listed in Table 1. We take the median map depth, in units of , for each of the observing elevations, and calculate their weighted mean based on the observing fraction at each elevation. Second, we estimate our sensitivities to point sources in the final, full-survey maps by spreading the array sensitivity accumulated over the survey duration over the survey area, using and an observing efficiency of , as described above. For both single-observation and full-depth maps, we obtain point source sensitivities by converting the map depth in to mJy using the beam FWHMs listed in Table 1 and applying an inverse-variance weighting with an atmospheric noise spectrum with a power-law index of and for GHz, respectively. Here, the values are estimated based on ACT data.121212Note that Ade et al. (2019) also made use of a set of values for the noise model used in forecasts (which is used in the cosmological forecasting in this paper as well), but there the term would have been better called a reference multipole, , rather than a knee multipole, which is usually defined as the multipole where the red noise caused by the atmosphere reaches the same level as the white noise; Ade et al. (2019) instead defined a “red noise” amplitude, , to scale the reference multipole. The relationship between the two is: , where is the power-law index of the red noise. For the point source forecasts in this paper, our choices of and are independent of the model used in Ade et al. (2019), but are still chosen to be consistent with the atmospheric noise measured by ACT. For the full-depth maps, we include additional “noise” contributions from the CMB and the CIB, since these terms dominate over the atmospheric noise in this case (other background sources like unresolved AGN are ignored). We ignore these terms for the single-observation maps since they are subdominant. The resulting sensitivities for both the single-observation and full-survey cases are given in Table 3.